### v1

# Inconsistency thresholds revisited: The effect of the graph associated with incomplete pairwise comparisons

Abstract
The inconsistency of pairwise comparisons remains difficult to interpret in the absence of acceptability thresholds. The popular 10% cut-off rule proposed by Saaty has recently been applied to incomplete pairwise comparison matrices, which contain some unknown comparisons. This paper revises these inconsistency thresholds: we uncover that they depend not only on the size of the matrix and the number of missing entries, but also on the undirected graph whose edges represent the known pairwise comparisons. Therefore, using our exact thresholds is especially important if the filling in patterns coincide for a large number of matrices, as has been recommended in the literature. The strong association between the new threshold values and the spectral radius of the representing graph is also demonstrated. Our results can be integrated into software to continuously monitor inconsistency during the collection of pairwise comparisons and immediately detect potential errors.
Keywords: Decision analysis; graph theory; inconsistency; incomplete pairwise comparisons; spectral radius
MSC class: 90B50, 91B08
JEL classification number: C44, D71
The mathematical structure known as a graph has the valuable feature of helping us to visualize, to analyze, to generalize, a situation of a problem we may encounter and, in many cases, assisting us to understand it better and possibly find a solution.111 Source: Benjamin et al., (2017, p. 1).
(Arthur Benjamin, Gary Chartrand, Ping Zhang: The Fascinating World of Graph Theory)
[…]
1 Introduction
Pairwise comparison matrices play a central role in several multi-criteria decision-making methodologies such as the Analytic Hierarchy Process (AHP) (Saaty,, 1977, 1980). However, while pairwise comparisons efficiently decompose complex decision-making problems into simple subproblems, this comes at a price: there is no guarantee of cardinal consistency. For example, if alternative $A$ is two times better than alternative $B$, and alternative $B$ is three times better than alternative $C$, alternative $A$ is not necessarily six times better than alternative $C$, in contrast to their indirect comparison through alternative $B$. Indeed, both empirical (Bozóki et al.,, 2013) and randomly generated (Csató and Petróczy,, 2021; Csató,, 2024) pairwise comparison matrices are usually inconsistent. The level of inconsistency is measured by inconsistency indices (Brunelli,, 2018).
1.1 The need for inconsistency thresholds in the incomplete case
The numerical value of inconsistency is difficult to interpret on its own, without a sharp threshold that separates acceptable and unacceptable levels of inconsistency. In the latter case, the original comparisons should be revised, for which several inconsistency reduction methods have been suggested (Mazurek et al.,, 2021). For instance, Bozóki et al., (2015) formulate nonlinear mixed-integer optimisation problems to determine the minimal number of matrix elements to be changed in order to reduce the inconsistency below a given value.
The first inconsistency index and threshold of acceptability have been proposed by the founder of the AHP methodology (Saaty,, 1977). It is based on dividing the inconsistency of the matrix by the average inconsistency of random pairwise comparison matrices, the so-called random index. Inconsistency can be tolerated if this ratio remains below 0.1, which directly provides a threshold for inconsistency.
Pairwise comparison matrices have a natural generalisation by allowing for missing entries (Harker,, 1987). Incomplete pairwise comparison matrices extend the range of potential applications to hundreds, or even millions of alternatives. For example, the pairwise comparisons of contestants in sports are usually intransitive (van Ours,, 2024, 2025) and may be easily incomplete if some pairs of contestants do not play against each other. Such case studies are presented by Bozóki et al., (2016); Chao et al., (2018); Csató, (2013); Petróczy and Csató, (2021); Temesi et al., (2024), among others. Analogously, neither completeness nor consistency holds for bilateral remittances (Petróczy,, 2021) or students’ preferences (Csató and Tóth,, 2020)
Even though some inconsistency indices are available for incomplete pairwise comparison matrices (Kułakowski and Talaga,, 2020; Szybowski et al.,, 2020), determining the associated thresholds is far from trivial. Naturally, an incomplete pairwise comparison matrix can be completed (Ágoston and Csató,, 2024; Bozóki et al.,, 2010; Csató et al.,, 2024; Tekile et al.,, 2023), but the completion methods are usually based on minimising inconsistency. Consequently, the thresholds derived for complete matrices become too permissive for incomplete matrices: they will accept the inconsistency of several incomplete matrices, which would be rejected unless the decision-maker provides the (almost) optimal values for all missing entries.
1.2 The research question and our contribution
According to our knowledge, there exists only one proposal of inconsistency thresholds for incomplete pairwise comparison matrices. Ágoston and Csató, (2022) adopt the idea of Saaty to compute the appropriate thresholds as a function of the number of alternatives and the number of missing entries. Unsurprisingly, the random index decreases if the number of unknown comparisons is higher because the inconsistency minimisation problem contains more variables, which generally implies a lower optimum.
Incomplete pairwise comparisons are often represented by undirected graphs where the vertices are the alternatives and the edges correspond to the known comparisons. The random indices, thus, the inconsistency thresholds depend on the number of vertices and edges (Ágoston and Csató,, 2022). However, is it not possible that other properties of the representing graph should also be considered?
The current paper aims to address this issue. Following the methodology of Ágoston and Csató, (2022), we compute the average inconsistency of a large number of random incomplete pairwise comparison matrices where not only the number of vertices and edges, but the graph itself is fixed, up to isomorphism. The inconsistency threshold turns out to be strongly related to the spectral radius of the graph. Furthermore, the structure of the edges influences the proportion of randomly generated matrices that have an acceptable level of inconsistency.
1.3 A motivating example
Consider the following incomplete pairwise comparison matrix:
[…]
Matrix $\mathbf{A}$ is inconsistent as $a_{12}\cdot a_{23}\cdot a_{34}=2\cdot 4\cdot 2=16$, but $a_{14}=5$. The dominant eigenvalue of the optimally completed matrix is $\lambda\approx 4.084$, resulting in an inconsistency index $\mathit{CI}\approx 0.0284$.
According to Ágoston and Csató, (2022, Table 2), the random index $\mathit{RI}\approx 0.306$ if there are $n=4$ alternatives and $m=2$ missing elements. However, Table 1 reveal that the random index is $\mathit{RI}\approx 0.265$ if the positions of the two missing entries (they are placed in different columns and rows) are taken into account, too.
Therefore, the inconsistency of matrix $\mathbf{A}$ can be tolerated based on Ágoston and Csató, (2022). However, this inconsistency is too high if the famous 10% rule of Saaty is applied in a more sophisticated way. For pairwise comparison matrices with a graph representation analogous to the above example, the novel threshold changes the decision on 1165 matrices whose entries are drawn from the Saaty scale.
1.4 Practical relevance
Recently, optimal filling in patterns have been recommended for incomplete pairwise comparison matrices that provide the closest weight vectors on average to the complete case (Szádoczki et al.,, 2022, 2023; Szádoczki and Bozóki,, 2025). When the pairwise comparisons are asked from the decision-makers in such a sequence, a large number of incomplete pairwise comparison matrices will have the same graph representation, see the experiment of Szádoczki et al., (2025). Then it is crucial not to use the naïve thresholds given by Ágoston and Csató, (2022) since they assume random positions for the missing entries in the matrix. The exact thresholds provided in the current paper apply to all incomplete pairwise comparison matrices with the same graph representation; hence, the same value remains valid for matrices collected from different decision-makers, which does not complicate the analysis of inconsistency but makes it more accurate.
Our results are also important for inconsistency monitoring (Bozóki et al.,, 2011). Since the pairwise comparison matrices are always filled sequentially by the decision-makers, we have a new incomplete pairwise comparison matrix after each step. By continuously checking the inconsistency of these matrices, any serious violation of inconsistency can be immediately reported to the decision-maker. Therefore, it is possible to instantly correct any error or misprint, even before all pairwise comparisons are collected. This might be especially useful due to the mental and time constraints of the experts.
1.5 Structure
The paper is organised as follows. The methodology is described in Section 2. Our main findings are presented and discussed in Section 3, while Section 4 offers concluding remarks.
2 Methodology
Inconsistency and incomplete pairwise comparison matrices are introduced in Section 2.1. Section 2.2 presents the background of computing the random index, and Section 2.3 gives an overview of our numerical analysis.
2.1 Inconsistency and incomplete pairwise comparison matrices
Let $n$ denote the number of alternatives. The pairwise comparisons of the alternatives are collected into a pairwise comparison matrix $\mathbf{A}=\left[a_{ij}\right]$, $a_{ij}>0$ means that the $i$th alternative is $a_{ij}$ times more preferred to the $j$th alternative. The matrix satisfies the reciprocity property: $a_{ji}=1/a_{ij}$ for all $1\leq i,j\leq n$.
A pairwise comparison matrix $\mathbf{A}$ is consistent if and only if $a_{ik}=a_{ij}\cdot a_{jk}$ holds for all $1\leq i,j,k\leq n$. In other words, the pairwise comparisons are consistent if the direct comparison of any two alternatives coincides with their indirect comparison through any third alternative. Otherwise, the matrix is called inconsistent.
According to the Perron–Frobenius theorem (Perron,, 1907; Frobenius,, 1908, 1909, 1912), a pairwise comparison matrix $\mathbf{A}$ has a dominant real eigenvalue $\lambda_{\max}\left(\mathbf{A}\right)$. The inconsistency index of Thomas L. Saaty (Saaty,, 1977) is:
$\mathit{CI}\left(\mathbf{A}\right)=\frac{\lambda_{\max}\left(\mathbf{A}\right)-n}{n-1}.$
$\mathit{CI}\left(\mathbf{A}\right)=0$ if and only if the pairwise comparisons are consistent.
Saaty suggests comparing $\mathit{CI}\left(\mathbf{A}\right)$ to the random index $\mathit{RI}$, the average $\mathit{CI}$ of a large number of pairwise comparison matrices whose entries above the diagonal are drawn independently and uniformly from the following Saaty scale:
$\left\{1/9,\,1/8,\,1/7,\dots,\,1/2,\,1,\,2,\dots,\,8,\,9\right\}.$
The values of the random index were computed several times as a function of the number of alternatives $n$ (Aguarón and Moreno-Jiménez,, 2003; Alonso and Lamata,, 2006; Bozóki and Rapcsák,, 2008; Csató and Petróczy,, 2021; Ozdemir,, 2005; Tummala and Ling,, 1998).
The inconsistency index divided by the random index defines the inconsistency ratio:
$\mathit{CR}\left(\mathbf{A}\right)=\frac{\mathit{CI}\left(\mathbf{A}\right)}{\mathit{RI}}.$
Incomplete pairwise comparison matrices allow for some missing comparisons. Therefore, $\mathbf{A}=\left[a_{ij}\right]$ is an incomplete pairwise comparison matrix if $a_{ij}>0$ or $a_{ij}=\star$ such that $a_{ij}>0$ implies $a_{ji}=1/a_{ij}$ and $a_{ij}=\star$ implies $a_{ji}=\star$. The number of missing comparisons above the diagonal is denoted by $m$, hence, $\lvert(i,j):a_{ij}=\star\rvert=2m$. Incomplete pairwise comparisons have a natural graph representation. Let $\mathbf{A}$ be an incomplete pairwise comparison matrix. The associated undirected graph is $G=(V,E)$, where the vertices correspond to the alternatives and $e_{ij}\in E$ is an edge if and only if the pairwise comparison $a_{ij}$ is known.
The adjacency matrix $\mathbf{B}=\left[b_{ij}\right]$ of graph $G=(V,E)$ is given by $b_{ij}=1$ if $e_{ij}\in E$ and $b_{ij}=0$ if $e_{ij}\notin E$. The spectral radius of a finite graph $G=(V,E)$ is the maximum of the absolute values of the eigenvalues of its adjacency matrix $\mathbf{B}$.
11223344
Figure 1: The graph representation of the pairwise comparison matrix $\mathbf{A}$ in Example 1
Example 1.
Take the following incomplete pairwise comparison matrix:
[…]
Its graph representation is given in Figure 1. The spectral radius of this graph equals 2.
Understanding the properties of the spectral radius is an active research field in discrete mathematics (Stevanović,, 2018). Denote the maximum and minimum degree of a graph $G$ by $\Delta(G)$ and $\delta(G)$, respectively. Hong, (1993) asked the following question more than three decades ago: If $G$ has the smallest possible spectral radius among all graphs with $n$ vertices and $e$ edges, does $\Delta(G)-\delta(G)\leq 1$ hold, namely, is $G$ almost regular? The answer seems to be positive, at least for dense graphs with $e\geq(n-1)(n-2)/2-2$ (Cioaba et al.,, 2024). Analogously, there are recent results on upper bounds of the spectral radius (Guo et al.,, 2019; Jin et al.,, 2024).
2.2 The eigenvalue minimisation problem and its solution
In order to complete an incomplete pairwise comparison matrix $\mathbf{A}$, Shiraishi and Obata, (2002) and Shiraishi et al., (1998) have proposed substituting the $m$ missing entries above the diagonal with positive variables collected in the vector $\mathbf{x}=\left[x_{k}\right]$, $1\leq k\leq m$. Since inconsistency ($\mathit{CI}$) is a monotone function of the dominant eigenvalue, the following optimisation problem is worth considering:
$\min_{\mathbf{x}>0}\lambda_{\max}\left(\mathbf{A}\left(\mathbf{x}\right)\right).$ (1)
(1) can be transformed into a convex optimisation problem (Bozóki et al.,, 2010, Section 3). The necessary and sufficient condition for the existence of a unique solution is the connectedness of the undirected graph (Bozóki et al.,, 2010, Theorem 2). Therefore, we will focus on incomplete pairwise comparison matrices represented by a connected graph. Bozóki et al., (2010, Section 5) presents an algorithm to obtain the optimal solution to (1).
As discussed by Ágoston and Csató, (2022), (at least) three possible approaches exist to calculate the random index for incomplete pairwise comparison matrices, depending on how the missing comparisons are related to the Saaty scale:
  • Method 1: all positive variables $x_{k}$ are unconstrained, the missing comparisons can be arbitrarily small or high for all $1\leq k\leq m$;
  • Method 2: all positive variables $x_{k}$ are constrained by the bounds of the Saaty scale, $1/9\leq x_{k}\leq 9$ for all $1\leq k\leq m$;
  • Method 3: all positive variables $x_{k}$ are constrained directly by the Saaty scale, $x_{k}$ should be either an integer from 1 to 9 or its reciprocal for all $1\leq k\leq m$.
Similar to Ágoston and Csató, (2022), Method 2 will be implemented to avoid numerical difficulties, as well as the problems caused by using a discrete scale for the variables in (1).
2.3 The computation approach
In contrast to Ágoston and Csató, (2022), the positions of the missing entries are not chosen randomly for a given number of unknown comparisons $m$, but the associated graph is fixed in advance. This does not make any difference if $m=1$ because all graphs with $n$ vertices and n(n1)/21n(n-1)/2-1 edges are isomorphic. However, it can be important if $m\geq 2$.
Consequently, the random index is computed as follows:
  1. 1.
    We choose a connected graph $G=(V,E)$ with $n$ vertices and $m$ missing edges compared to a complete graph;
  2. 2.
    We generate an incomplete pairwise comparison matrix $\mathbf{A}$ with graph representation $G$ such that each element of the Saaty scale is drawn with a probability of $1/17$ for each known comparison above the diagonal;
  3. 3.
    We solve the minimisation problem (1) by restricting all variables in $\mathbf{x}>0$ between $1/9$ and 99 (Method 2 in Section 2.2), which gives a complete pairwise comparison matrix $\mathbf{\hat{A}}$ as the optimal solution;
  4. 4.
[…]
  5. 5.
    We repeat Steps 24 to get 1 million random matrices with a given graph representation $G$;
  6. 6.
    We compute the mean of the inconsistency indices $\mathit{CI}$ from Step 4.
3 Results
Section 3.1 provides a comprehensive analysis in the case of four alternatives, and Section 3.2 focuses on pairwise comparison matrices with two missing entries. Finally, matrices of size five and six are discussed in Section 3.3.
3.1 Incomplete matrices of size four
If $n=4$, only the case $m=2$ unknown entries is interesting. For $m=1$, only one graph exists since $a_{14}=\star$ can be assumed without loss of generality. For $m=3$, a connected graph is necessarily a spanning tree, and the incomplete pairwise comparison matrix has a consistent completion if the missing entries remain unconstrained.
If $m=2$, there are two possible graphs: (1) the two missing edges are independent, they do not share any vertex (Figure 1); (2) the two missing edges have a common vertex. It is easy to check that the probability of the first case (1) equals
$\frac{n(n-1)/2-(n-1+n-2)}{n(n-1)/2-1}=1-\frac{4n-8}{n(n-1)-2},$
which is only $1/5=20\%$ for $n=4$, but converges to $1$ when $n$ increases.
The value of the random index $\mathit{RI}$ can be computed exactly: the number of possible scenarios is $17^{4}=83{,}521$ for both graphs due to the four known entries and the 17 elements of the Saaty scale. Hence, in contrast to what is done in Ágoston and Csató, (2022), we carry out a complete enumeration instead of generating 1 million random matrices.
Table 1: Random indices and their implications for
incomplete pairwise comparison matrices of size four
a Naïve random index $\mathit{RI}$ that is independent of the graph (Ágoston and Csató,, 2022)
Graph 1 Graph 2
Random index 0.3061 0.3061
Acceptable matrices ($\mathit{CI}\leq 0.1$) 14,789 12,095
Unacceptable matrices ($\mathit{CI}>0.1$) 68,732 71,426
Ratio of acceptable matrices 17.71% 14.48%
b Sophisticated random index $\mathit{RI}$ that depends on the graph (this paper)
Graph 1 Graph 2
Random index 0.2646 0.3165
Acceptable matrices ($\mathit{CI}\leq 0.1$) 13,633 12,343
Unacceptable matrices ($\mathit{CI}>0.1$) 69,888 71,178
Ratio of acceptable matrices 16.32% 14.78%
Table 1 reports the values of the refined random indices and their implications. The structure of the graph has a non-negligible effect; $\mathit{RI}$ for Graph 1 is smaller by more than 15% compared to $\mathit{RI}$ for Graph 2. Ignoring the graph misclassifies 1156 and 248 incomplete pairwise comparison matrices in the two possible cases, respectively. Even though the standard $\mathit{CI}$ threshold of 0.1 becomes more (less) restrictive for Graph 1 (Graph 2) if the structure of the graph is taken into account, the number of matrices with an acceptable inconsistency is still higher for Graph 2 than for Graph 1.
3.2 Incomplete matrices with two missing elements
[…]
Figure 2: Spectral radius and random index for incomplete
pairwise comparison matrices with $m=2$ missing entries
Figure 2 shows the random indices for the two different incomplete pairwise comparison matrices with two missing entries as a function of the spectral radius of the representing graph if the number of alternatives is between four and nine. The plot has two important messages. First, a higher spectral radius is always associated with a higher value of $\mathit{RI}$ if $m=2$. Second, the difference between the two cases gradually disappears when the number of alternatives grows, that is, the random indices provided in Ágoston and Csató, (2022) become less distorted. This is not surprising since fixing the number of unknown comparisons means that the ratio of known elements increases together with $n$.
3.3 Incomplete matrices of size five and six
a $G_{2,1}:\left[4,4,3,3,2\right]$
b $G_{2,2}:\left[4,3,3,3,3\right]$
c $G_{3,1}:\left[4,3,3,3,1\right]$
d $G_{3,2}:\left[4,3,3,2,2\right]$
e $G_{3,3}:\left[3,3,3,3,2\right]$
f $G_{3,4}:\left[4,4,2,2,2\right]$
g $G_{4,1}:\left[3,3,2,2,2\right]$
h $G_{4,2}:\left[3,3,2,2,2\right]$
i $G_{4,3}:\left[4,2,2,2,2\right]$
j $G_{4,4}:\left[4,3,2,2,1\right]$
k $G_{4,5}:\left[3,3,3,2,1\right]$
l $G_{5,1}:\left[4,2,2,1,1\right]$
m $G_{5,2}:\left[3,3,2,1,1\right]$
n $G_{5,3}:\left[3,2,2,2,1\right]$
o $G_{5,4}:\left[3,2,2,2,1\right]$
p $G_{5,5}:\left[2,2,2,2,2\right]$
Figure 3: Possible graph representations of incomplete pairwise comparison matrices of size five with at least two and at most five missing entries
Notes: Dotted lines indicate the known comparisons, thick red lines indicate the unknown comparisons.
Graph $G_{i,j}$ has $i$ missing edges. The vectors show the degree distribution of the associated graph.
See Table A.1 for the spectral radii and random indices of these incomplete pairwise comparison matrices.
If $n=5$, all cases are investigated, that is, the number of missing entries varies from 1 to 5 because $m=6$ implies the existence of a consistent completion (if the unknown entries remain unconstrained). Figure 3 presents the 16 possible graphs if $n=5$ and $2\leq m\leq 5$.
Hence, there are $1+2+4+5+5=17$ random indices, depending on the structure of the graph. All of them are reported in Table A.1 in the Appendix, together with their probability of occurrence if the number of missing entries $m$ is fixed. The probabilities are estimated by simulations, contrary to the exact values derived in Section 3.1. For a fixed value of $m$, the spectral radius is the smallest if the graph is (almost) regular ($G_{2,2}$, $G_{3,3}$, $G_{4,1}$, $G_{5,5}$). These graphs have the smallest probability to occur except for $G_{3,3}$, and the corresponding random index $\mathit{RI}$ is lower than for any other graphs having the same number of edges.
a $G_{1}:\left[4,4,3,3,2,2\right]$
b $G_{2}:\left[4,4,3,3,2,2\right]$
c $G_{3}:\left[4,3,3,3,3,2\right]$
d $G_{4}:\left[5,4,3,2,2,2\right]$
e $G_{5}:\left[4,4,3,3,2,2\right]$
f $G_{6}:\left[5,3,3,3,3,1\right]$
g $G_{7}:\left[5,3,3,3,2,2\right]$
h $G_{8}:\left[4,3,3,3,3,2\right]$
i $G_{9}:\left[5,4,3,3,2,1\right]$
j $G_{10}:\left[4,4,4,2,2,2\right]$
k $G_{11}:\left[4,4,4,3,2,1\right]$
l $G_{12}:\left[5,3,3,3,2,2\right]$
m $G_{13}:\left[4,4,3,3,3,1\right]$
n $G_{14}:\left[4,4,3,3,3,1\right]$
o $G_{15}:\left[4,4,3,3,2,2\right]$
p $G_{16}:\left[4,3,3,3,3,2\right]$
q $G_{17}:\left[4,4,3,3,2,2\right]$
r $G_{18}:\left[3,3,3,3,3,3\right]$
s $G_{19}:\left[5,5,2,2,2,2\right]$
t $G_{20}:\left[3,3,3,3,3,3\right]$
Figure 4: Possible graph representations of incomplete pairwise comparison matrices of size six with six missing entries
Notes: Dotted lines indicate the known comparisons, thick red lines indicate the unknown comparisons.
The vectors show the degree distribution of the associated graph. See Table Appendix for the spectral radii and random indices of these incomplete pairwise comparison matrices.
If $n=6$, we have computed the random indices only up to $m=7$ as the number of different graphs increases with $m$; for instance, 20 scenarios exist for $n=6$ and $m=6$ according to Figure 4. The random indices are provided in Table Appendix in the Appendix. Again, the regular graph $G_{20}$ in Figure 4 occurs with the smallest probability and has the smallest value of $\mathit{RI}$ if $m=6$, and the other regular graph $G_{18}$ has the second smallest random index.
a $n=5$ alternatives
[…]
b $n=6$ alternatives
[…]
Figure 5: Spectral radius and random index for incomplete matrices of size five and six
Figure 5 presents the random indices for incomplete pairwise comparison matrices of size five (Figure 5.a) and six (Figure 5.b), as a function of the spectral radius of the associated graph. The monotonic relationship between the spectral radius and $\mathit{RI}$ breaks first when $n=5$ and $m=4$ (see also Table A.1). However, a higher spectral radius usually implies a higher random index even for $n=5$ and $n=6$. If the number of alternatives is six, the spectral radii can overlap for different values of missing entries $m$, but the random indices still monotonically decrease as a function of $m$. In other words, the number of variables in the optimisation problem (1) is more important than their relative position in the representing graph with respect to the optimum, which is intuitive.
Table 2: Extreme random indices for incomplete pairwise comparison
matrices of a fixed size $n$ and number of missing entries $m$
$n$ $m$ Minimal $\mathit{RI}$ Maximal $\mathit{RI}$ Ratio
4 2 0.2646 0.3165 83.61%
5 2 0.7275 0.7454 97.60%
5 3 0.5377 0.5926 90.73%
5 4 0.3426 0.3952 86.70%
5 5 0.1717 0.2256 76.13%
6 2 1.0007 1.0099 99.09%
6 3 0.8658 0.8983 96.38%
6 4 0.7431 0.8059 92.20%
6 5 0.6126 0.6702 91.40%
6 6 0.4733 0.5402 87.61%
6 7 0.3562 0.4312 82.60%
Table 2 shows the extreme values of the random index $\mathit{RI}$ for given pairs of parameters $m$ and $n$. Unsurprisingly, the influence of the graph decreases if $m$ is fixed and the number of alternatives grows, but increases if $n$ is fixed and the number of missing entries increases. Among the 11 scenarios, the random index can be smaller by more than 10% in five cases, and the reduction may exceed 23% if both parameters are equal to five. This reinforces that the underlying graph cannot be ignored in the calculation of inconsistency thresholds for incomplete pairwise comparison matrices.
a $n=5$ alternatives
[…]
b $n=6$ alternatives
[…]
Figure 6: Random index and the probability of acceptable
inconsistency for incomplete matrices of size five and six
Finally, Figure 6 presents the relationship between the random index and the proportion of randomly generated incomplete pairwise comparison matrices with the same graph representation that have an acceptable inconsistency. The effect of the graph becomes more important if the number of missing entries $m$ is higher. Interestingly, a smaller (more restrictive) random index may lead to a higher probability of acceptable inconsistency, similar to what can be seen in Table 1 for $n=4$ alternatives.
4 Conclusions
Our paper provides refined random indices for incomplete pairwise comparison matrices in order to determine the threshold of acceptability for their inconsistency. The values of Ágoston and Csató, (2022) are made more accurate by considering not only the number of alternatives and missing entries, but also the graph representing the known comparisons. We show that the graph structure has a non-negligible effect and the random index is in a strict—albeit not perfect—relationship with the spectral radius of the corresponding graph. Furthermore, the underlying graph influences the ratio of randomly generated pairwise comparison matrices whose inconsistency can be accepted, which might question the validity of the 10% rule of Saaty.
The research is far from closed. First, the random indices are fully computed only up to five alternatives due to the complexity of the numerical calculations. Second, even though the association with the spectral radius is surprisingly strong, other measures of graph “robustness” may provide an even better fit. The association between the random index and the spectral radius may also turn out to be a mere correlation, which is explained by a third, currently hidden indicator.
Acknowledgements
We acknowledge the Digital Government Development and Project Management Ltd. for awarding us access to the Komondor HPC facility based in Hungary.
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838 and the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
Appendix
Table A.1: Random indices for incomplete pairwise comparison matrices of size $n=5$
Value of $m$ Graph Probability Spectral radius Random index $\mathit{RI}$ Ratio of acc.
1 1 100.00% 3.6458 0.9246
2 1 67.05% 3.3234 0.7454 1.16%
2 2 32.95% 3.2361 0.7275 1.19%
3 1 16.92% 3.0861 0.5926 2.43%
3 2 49.49% 2.9354 0.5535 2.46%
3 3 25.28% 2.8558 0.5377 2.58%
3 4 8.31% 3 0.5611 2.44%
4 1 4.62% 2.4495 0.3426 5.46%
4 2 29.59% 2.4812 0.3557 5.29%
4 3 6.89% 2.5616 0.3745 4.99%
4 4 29.05% 2.6855 0.3927 4.98%
4 5 29.85% 2.6412 0.3952 5.18%
5 1 12.92% 2.3429 0.2190 10.19%
5 2 27.28% 2.3028 0.2235 10.63%
5 3 27.31% 2.1358 0.1899 11.01%
5 4 26.76% 2.2143 0.2256 11.13%
5 5 5.72% 2 0.1717 11.15%
  • Notes: The column Probability shows the probability that the given graph occurs if all missing comparisons are placed randomly for a fixed pair of parameters $m$ and $n$.
  • The column Ratio of acc. shows the probability that an incomplete pairwise comparison matrix with the given graph representation, whose known entries are drawn uniformly from the Saaty scale, has an acceptable level of inconsistency, for a fixed pair of parameters $m$ and $n$.
Random indices for incomplete pairwise comparison matrices of size $n=6$
Value of $m$ Graph Probability Spectral radius Random index $\mathit{RI}$ Ratio of acc.
[…]
Value of $m$ Graph Probability Spectral radius Random index $\mathit{RI}$ Ratio of acc.
[…]
2 1 56.87% 4.4279 1.0099 0.05%
2 2 43.13% 4.3723 1.0007 0.04%
3 1 13.16% 4.2015 0.8983 0.10%
3 2 39.95% 4.1190 0.8841 0.10%
3 3 3.34% 4 0.8658 0.10%
3 4 4.39% 4.1623 0.8904 0.10%
3 5 39.16% 4.0678 0.8765 0.10%
4 1 2.10% 4.0514 0.8059 0.21%
4 2 6.60% 3.7321 0.7457 0.21%
4 3 4.50% 3.7664 0.7498 0.22%
4 4 26.75% 3.8590 0.7659 0.21%
4 5 25.76% 3.7785 0.7525 0.21%
4 6 13.58% 3.7136 0.7431 0.21%
4 7 4.31% 3.8201 0.7597 0.21%
4 8 13.12% 3.8951 0.7700 0.21%
4 9 3.28% 3.8284 0.7600 0.20%
5 1 12.48% 3.4679 0.6267 0.44%
5 2 12.35% 3.5344 0.6356 0.43%
5 3 11.74% 3.5926 0.6429 0.44%
5 4 12.15% 3.4979 0.6297 0.46%
5 5 6.00% 3.7105 0.6694 0.44%
5 6 11.81% 3.5141 0.6310 0.45%
5 7 2.07% 3.4495 0.6219 0.44%
5 8 11.80% 3.3885 0.6141 0.45%
5 9 3.17% 3.6262 0.6464 0.43%
5 10 6.05% 3.4609 0.6235 0.46%
5 11 3.92% 3.6903 0.6702 0.44%
5 12 1.51% 3.3723 0.6126 0.44%
5 13 3.00% 3.5616 0.6412 0.44%
5 14 1.95% 3.3923 0.6127 0.46%
6 1 7.38% 3.1819 0.5055 0.89%
6 2 3.92% 3.2361 0.5111 0.93%
6 3 7.48% 3.0868 0.4915 0.90%
6 4 7.24% 3.2814 0.5130 0.90%
6 5 14.59% 3.1692 0.4994 0.94%
6 6 2.01% 3.3234 0.5259 0.91%
6 7 6.87% 3.2227 0.5069 0.89%
6 8 6.92% 3.1149 0.4933 0.94%
6 9 7.18% 3.4037 0.5389 0.87%
6 10 2.36% 3.2361 0.5097 0.89%
6 11 7.14% 3.3839 0.5393 0.91%
6 12 1.44% 3.2618 0.5169 0.88%
6 13 3.28% 3.3539 0.5402 0.93%
6 14 7.54% 3.2948 0.5267 0.92%
6 15 1.93% 3.1413 0.4940 0.95%
6 16 7.34% 3.0922 0.4882 0.97%
6 17 3.64% 3.1888 0.5023 0.93%
6 18 1.23% 3 0.4782 0.95%
6 19 0.32% 3.3723 0.5222 0.89%
6 20 0.16% 3 0.4733 1.03%
7 1 10.93% 2.9809 0.4030 1.87%
7 2 11.70% 3.0143 0.4016 1.81%
7 3 2.87% 3.1642 0.4293 1.82%
7 4 2.92% 2.8951 0.3920 1.91%
7 5 6.32% 2.9439 0.3909 1.88%
7 6 11.43% 2.8529 0.3760 1.89%
7 7 2.95% 2.7913 0.3707 1.90%
7 8 5.84% 2.8136 0.3706 1.85%
7 9 2.10% 3.1020 0.4056 1.79%
7 10 5.95% 2.9327 0.3903 1.88%
7 11 3.07% 2.9032 0.3783 1.86%
7 12 1.88% 3.0965 0.4312 1.94%
7 13 2.87% 3.0478 0.4082 1.88%
7 14 5.59% 3.0437 0.4003 1.78%
7 15 2.83% 2.8422 0.3847 1.80%
7 16 6.34% 2.7964 0.3648 1.92%
7 17 2.82% 2.7321 0.3682 1.85%
7 18 0.87% 3.1774 0.4273 1.77%
7 19 5.76% 2.7411 0.3598 1.95%
7 20 1.62% 2.7321 0.3562 2.00%
7 21 0.28% 2.8284 0.3633 1.94%
7 22 3.07% 2.9474 0.3864 1.75%